OpenStudy (swagster23780):

Multiply. Express your answer in lowest terms. 1/4 .3/5 =

OpenStudy (haleyelizabeth2017):

\[\frac{1}{4} \times\frac{3}{5}\] yes?

OpenStudy (swagster23780):

Yes

OpenStudy (haleyelizabeth2017):

\[\frac{1}{4} \times \frac{3}{5}=\frac{1\times3}{4\times5}=?\]

OpenStudy (anonymous):

3/20

OpenStudy (haleyelizabeth2017):

Please no direct answers @ethan726 :) thank you

OpenStudy (swagster23780):

Is It Correct

OpenStudy (haleyelizabeth2017):

Yes, but would you have known that without him saying that?

OpenStudy (swagster23780):

Prob

OpenStudy (swagster23780):

ably